RC Meeting and Runplan, Friday, Sept 28
- Yesterday the HMS quad work was completed by about 1pm and we started the optics study of HMS/SHMS at high momentum.
- Beam was bad until around 6pm when MCC had solved the problem of the many trips coming from Hall B beam.
- When changing to set the HMS at 6 GeV , The HMS Q3 would not stay at a given set current. This put into doubt the the higher momentum HMS optics data (taken at 6.6 and 6.3).
- MCC was able to deliver up to 60uA on the owl shift. Given problem with HMS, decided to do the BCM which was completed. https://logbooks.jlab.org/entry/3599568
- After BCM data was taken , then continued with the SHMS optics which was completed.
- Turned on EDTM pulser for BCM calib data. Left it on.
- The Hall went into controlled access around 8:30 for work on the HMS Q3. Found that the opto-isolator was not pushed in all the way.
- Around 9:00, MCC started the pass change for Hall A.
- Around noon, started to receive beam and restarted the optics study for HMS ( repeated some SHMS). finsihed around 4:20.
- Updated the EPICS BCM names and calibrations https://logbooks.jlab.org/entry/3599765
- Start heepcheck runs ( about 9 hours).
- Steve Wood to comment on importance of watching closely the GUI and TV screen when moving spectrometers.
- Changing the attenuator of the SHMS Cer sum that is used to monitor that leg of the trigger for 0.3 to 1.0 . https://logbooks.jlab.org/entry/3599768
- HMS UPS is bad. Jumpered interlock. https://logbooks.jlab.org/entry/3599763
- Gas shed https://logbooks.jlab.org/entry/3599543